OpenStudy (anonymous):

integrate 1/(x-2)(x+4) when x>2

OpenStudy (anonymous):

seoarate fractions: \[\frac{A}{x-2}+\frac{B}{x+4}=\frac{A(x+4)+B(x-2)}{(x-2)(x+4)}=\frac{Ax+4A+Bx-2B}{(x-2)(x+4)}\] so A+B=0 and 4A-2B=1 solving this system you faind that A=1/6 and B=-1/6. So your integral becomes: \[\frac{1}{6}\int\limits\frac{dx}{(x-2)}-\frac{1}{6}\int\limits\frac{dx}{x+4}=1/6Ln(x-2)-1/6Ln(x+4) +C\]

OpenStudy (anonymous):

It also needs to be evaluated on the interval (2,infinity)

OpenStudy (anonymous):

it's same, as long the (x-2)>0